### v4

# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

Abstract.
Quantization of foundational models (FMs) is significantly more challenging than traditional DNNs due to the emergence of large magnitude values called outliers. Existing outlier-aware algorithm-architecture co-design techniques either use mixed-precision, retaining outliers at high precision but compromise hardware efficiency, or quantize inliers and outliers at the same precision, improving hardware efficiency at the cost of accuracy. To address this mutual exclusivity, we propose MicroScopiQ, a novel co-design technique that leverages pruning to complement outlier-aware quantization. MicroScopiQ retains outliers at higher precision while pruning a certain fraction of least important weights to distribute the additional outlier bits; ensuring high accuracy, aligned memory and hardware efficiency. We design a high-throughput, low overhead accelerator architecture composed of multi-precision INT processing elements and a network-on-chip called ReCoN that efficiently abstracts the complexity of supporting high-precision outliers. Additionally, unlike prior techniques, MicroScopiQ does not assume any locality of outlier weights, enabling applicability to a broad range of FMs. Extensive experiments across diverse quantization settings demonstrate that MicroScopiQ achieves state-of-the-art quantization accuracy, while delivering up to $\mathbf{3}\times$ faster inference and $\mathbf{2}\times$ lower energy consumption compared to existing alternatives. Code is available at: MicroScopiQ-LLM-Quantization.git
Foundational Models, Quantization, Pruning, Hardware Accelerator
journalyear: 2025copyright: ccconference: Proceedings of the 52nd Annual International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanbooktitle: Proceedings of the 52nd Annual International Symposium on Computer Architecture (ISCA ’25), June 21–25, 2025, Tokyo, Japandoi: 10.1145/3695053.3730989isbn: 979-8-4007-1261-6/2025/06ccs: Computer systems organization Systolic arraysccs: Computer systems organization Neural networksccs: Computer systems organization Data flow architecturesccs: Networks NoC
1. Introduction
Recent advancements in AI (Goodfellow et al., 2014; Li et al., 2022; Touvron et al., 2023; Chen et al., 2024) have been propelled by a class of models called foundational models (FMs), which encompass large language models (LLMs) and vision-language models (VLMs). FMs leverage billion-scale parameters for improved learning (Yang et al., 2024; Zhang et al., 2024) but impose substantial demands on memory, energy, and compute resources. Recent research has focused on various model compression techniques such as pruning (Frantar and Alistarh, 2023; Ashkboos et al., 2024a; Yin et al., 2023) and quantization (Lin et al., 2024a; Shao et al., 2023; Ramachandran et al., 2024b) to reduce memory and computational overhead, enabling efficient FM inference on resource-constrained devices.
Model pruning (Frantar and Alistarh, 2023) reduces memory footprint by removing ineffectual model parameters, such as individual weights (unstructured) or blocks of weights (structured), and storing sparse tensors in a compressed format (Jeong et al., 2023). However, pruning of FMs may be infeasible due to, significant accuracy drops even at low pruning ratios (Kuzmin et al., 2024; Yin et al., 2024) and potential demand for compute and memory-intensive fine-tuning to regain accuracy. Model quantization, on the other hand, reduces the size of a target model by representing weights and/or activations at low precision (Ramachandran et al., 2024b; Kang et al., 2024; Ramachandran et al., 2024a; Kundu et al., 2025). Recent works on quantization (Dettmers et al., 2022; Guo et al., 2023; Ramachandran et al., 2025b) have identified that quantizing LLMs is considerably more challenging than quantizing traditional DNNs (Li et al., 2021; Tambe et al., 2020) due to the emergence of large magnitude features known as outliers (Zadeh et al., 2020). These outliers significantly impact model accuracy and require specialized handling (Guo et al., 2023) compared to inliers.
Table 1. MicroScopiQ vs. prior outlier-aware quantization techniques, categorized into two groups, A (Zadeh et al., 2020; Park et al., 2018), B (Guo et al., 2023).
  Methods
Categories Group A Group B MicroScopiQ
  Accuracy [Uncaptioned image]  High [Uncaptioned image]  Low [Uncaptioned image]High
  Effective bit-width [Uncaptioned image]  High (18.17b) [Uncaptioned image]  Low (2b) [Uncaptioned image]Low (2.36b)
  Flexibility [Uncaptioned image]  No [Uncaptioned image]  No [Uncaptioned image]Yes
  Aligned memory [Uncaptioned image]  Unaligned [Uncaptioned image]  Aligned [Uncaptioned image]Aligned
  PE design [Uncaptioned image]  Complex [Uncaptioned image]  Complex [Uncaptioned image]Simple
  HW overhead [Uncaptioned image]  High [Uncaptioned image]  Moderate [Uncaptioned image]Low
To address the issue of outliers in FMs, recent algorithm/ architecture co-design techniques (Shao et al., 2023; Lin et al., 2024a; Guo et al., 2023) have proposed different types of outlier-aware quantization. These techniques can be broadly categorized based on their outlier handling approach: A Maintaining outliers at higher precision compared to the inliers, or B Quantizing outliers at the same precision as inliers while using different data formats or scaling factors for outliers.
Techniques in group A, such as OWQ (Lee et al., 2024), SpQR (Dettmers et al., 2023), SDQ (Jeong et al., 2024) (algorithm) and GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018)(architecture co-design) exhibit low accuracy degradation. This is because, they typically store outliers at high precision separated from lower precision inliers. However, these techniques result in, (a) low compression factor with high effective bit-width (EBW [^1] The average number of bits used to represent each quantized parameter of a model.) and (b) inefficient hardware and unaligned memory access.
On the other hand, techniques in group B, such as AWQ (Lin et al., 2024a) (algorithm) and OliVe (Guo et al., 2023) (architecture co-design) quantize outliers at the same precision as inliers following different strategies. AWQ tries to identify a separate outlier-specific scale factor via channel-wise scaling. OliVe uses the “flint” data format (Guo et al., 2022) for inliers and “abfloat” (Guo et al., 2023) for outliers, both at 4-bit precision. These techniques mitigate the unaligned memory access while providing high compression. However, they suffer from significant accuracy degradation, particularly at ultra-low bit widths. This may be attributed to the reduced representational range available to outliers at ultra low-precision. Additionally, these methods (Guo et al., 2023) rely on a specific kind of locality of presence for outliers, that might not be true for all FMs, as we shall demonstrate in this work (section 3.1).
Based on the shortcomings of existing solutions discussed above, we identify that assignment of higher bit-width for outliers is essential for good accuracy while for aligned memory and hardware efficiency a consistent bit-budget and data type per tensor element is desired. Here, by consistency we mean that on average each scalar within a tensor should be represented by a fixed bit-width of a particular data-type. However, these demands are conflicting.
Contributions. To provide a unified solution, we investigate on a fundamental question:
[…]
Towards achieving this, we present a novel co-design technique for the post-training quantization (PTQ) of FMs, namely MicroScopiQ. Our approach effectively leverages pruning with outlier aware quantization to achieve both memory alignment and improved accuracy. To effectively perform this for a layer, we quantize outliers at twice the precision of inliers and prune the least important weights based on the Hessian information. We then redistribute the additional bits of the outlier weights in these pruned locations. This ensures memory alignment while allowing outlier weights to have higher precision. Additionally, to reduce error we use the recently proposed MicroScaling (MX) FP data format (Rouhani et al., 2023; Darvish Rouhani et al., 2023) to quantize outlier weights as opposed to MX-INT inlier quantization. While prior work such as SDQ (Jeong et al., 2024) also combined pruning and quantization, it contrasts with our approach in its limited outlier flexibility, lower compression factor, and unaligned memory access.
To efficiently support outliers in a different format with different bits at different locations in hardware, we present an intelligent NoC architecture called, Redistribution and Coordination NoC (ReCoN). It offers minimal overhead and high throughput outlier processing and reorganization. We then present an accelerator that leverages ReCoN with a simple, homogeneous INT-PE array. Additionally, we extend the accelerator to be generic enough to support multiple bit-precision (2/4-bit operations). As summarized in Table 1, MicroScopiQ blends the advantages of group A and B techniques while mitigating their specific drawbacks.
Our key contributions can be summarized as follows:
  • We present MicroScopiQ, a PTQ framework to efficiently integrate pruning with outlier-aware quantization (section 4).
  • To effectively deploy MicroScopiQ in a systolic array architecture we present a novel architecture supporting multi-precision, homogeneous PEs with a low-overhead NoC architecture (section 5).
[…]
2. Background
2.1. Model Quantization
A typical quantization (Ramachandran et al., 2024a; Li et al., 2021; Ramachandran et al., 2025b) process involves two steps: establishing quantization parameters given the quantization data format $(\tau)$ and bit-width $(b)$, and mapping the high-precision tensor to the quantized representation. For a typical symmetric quantization (Zhao et al., 2024) (zero-point is 0) of a tensor $X$, the scale factor ($s$) is given by,
(1) $s=\frac{\emph{max}(X)}{\emph{max}_{\tau}^{b}}$
$\emph{max}_{\tau}^{b}$ is the maximum representable value of a data format (Ramachandran et al., 2024a). For $b$-bit INT quantization, $\emph{max}_{INT}^{b}=2^{b-1}-1$. After determining the quantization parameters, the quantized tensor is given by (Ramachandran et al., 2024a),
(2) $Q(X,s,b)=\emph{clip}(\Bigl{\lfloor}{\frac{X}{s}}\Bigr{\rceil},\emph{min}_{\tau}^{b},\emph{max}_{\tau}^{b})$
In model quantization, the quantization parameters can be shared at different granularity for different accuracy-overhead trade-offs. In increasing order of overheads, we have per-tensor quantization, wherein the scale factor is shared among all tensor elements. In per-channel quantization, the scale factor is shared per row/column of a tensor. Finally, in group quantization, the parameters are shared at a finer granularity between groups of $k$ (64, 128 etc.) elements in a row or column. These groups are formed by dividing channels into multiple non-overlapping contiguous blocks. In this paper, we adopt MX-INT and MX-FP quantization for inliers and outliers, respectively.
Figure 1. Depiction of MX-FP data format with level-1 scale factor and level-2 microExponent ($\mu X$), with $k_{1}$ and $k_{2}$ the group sizes over which these two factors are shared.
2.2. Microscaling Data Format
The MX data format proposed by prior works (Rouhani et al., 2023; Darvish Rouhani et al., 2023), is standardized by the Open Compute Project (Project, 2024) with support from Microsoft, Intel, NVIDIA and others. As shown in Figure 1, MX is a variant of block data representation (BDR) (Darvish Rouhani et al., 2023) that defines a format to represent a group of values collectively using shared scale factors. It leverages multi-level, power-of-two scaling, at fine- (level-1, $k_{1}$) and ultra-fine (level-2, $k_{2}$) granularity (Dai et al., 2021; Drumond et al., 2018). The MX data format is characterized by four components: i) scale factors (level-1, 2), ii) data type ($\tau$), iii) bit-width ($b$) and iv) group sizes ($k_{1},k_{2}$). In this paper we denote an MX-FP format as MX-FP-$b$${}_{k_{1},k_{2}}$. In this work, we adopt the version of the MX-FP data format proposed in (Rouhani et al., 2023), employing multi-level scaling. The level-1 scale factor for MX-FP is computed following Equation 1. Conversely, for level-2 scale factor, we identify that MX-FP leverages the sharing of exponent field of FP values (Wang et al., 2018) (referred as $\mu X$ in Figure 1). We show in section 4.2 that by taking advantage of this insight i.e., the concept of shared $\mu X$, we are able to represent FP-outliers in INT format, thereby, enabling the design of simple, homogeneous INT-based PEs. For inliers, we employ MX-INT-$b$${}_{k_{1}}$ with a single level of scale factor following (Project, 2024). This is because, INT format does not possess an exponent field, thereby, a level-2 scale factor similar to MX-FP is not applicable. For simplified understanding, MX-INT-$b$${}_{k_{1}}$ inlier quantization can be viewed as analogous to INT group quantization utilizing an E8M0 scale factor.
Figure 2. (a) Layer-wise distribution of outliers and adjacent outliers as a percentage of total number of weights, (b) Quantization accuracy comparison between OliVe-W4A16 and MicroScopiQ-W2A16 on various benchmarks.
3. Motivation
3.1. Limitations of existing techniques
In Table 1, we compare candidate proposals from group A: GOBO (Zadeh et al., 2020) and group B: OliVe (Guo et al., 2023) across various metrics. GOBO is able to achieve high accuracy by retaining outliers at full-precision. It stores outliers separately from low-precision inliers by using sparse representations with the associated outlier indices (see Figure 3(b)). By retaining outliers at full-precision, GOBO results in high EBW. Moreover, the compressed sparse storage and multiple precisions results in unaligned and random memory accesses (Guo et al., 2023), significantly impacting inference latency. Furthermore, GOBO’s outlier handling is hardware inefficient, requiring complex PEs. Similarly, a recent work (Jeong et al., 2024) proposed to decompose a vector of weights into two separate inlier and outlier vectors each quantized in different precisions with outliers at a higher precision.
OliVe (Guo et al., 2023) proposes a scheme to ensure aligned memory access by quantizing inliers and outliers at the same precision (low EBW), but using different data formats. To enable differentiation between the inlier and outlier formats, it prunes the value adjacent to the outlier for use as an identifier (see Figure 3(c)). However, OliVe results in significant accuracy degradation, especially at low precision (see Figure 2(b)), due to: 1) sacrificing a number encoding from inliers for exclusive use as an identifier, reducing the number of representable values in the quantized range, and 2) the rigid assumption of outlier locality–that outliers are never adjacent to each other and only inliers are almost always adjacent to outliers (see section 3.2), leading to unintended outlier pruning. Furthermore, OliVe requires a fairly complex PE design incurring significant encoding/decoding overheads to convert the different formats into a unified processing format (exponent-integer pair). In this paper, we show that despite quantizing outliers at higher-precision and in a different format, we ensure aligned memory access, simple PE design and minimal hardware overhead.
3.2. Adjacent Outliers Matter
Similar to prior works (Guo et al., 2023; Park et al., 2018), we leverage the $3\sigma$ rule (Pukelsheim, 1994) to categorize weights as outliers. We visually demonstrate the distribution of outliers and adjacent outliers  [^2] We define adjacent outliers as two contiguous outliers along the dot-product dimension (see row 2 of the LLM weight matrix in the center of Figure 3). as a percentage of the total number of weights in a layer across different FMs in Figure 2(a). As the orange box-plot shows, outliers depict a maximum percentage of $\sim 5.1\%$. Outliers are prevalent in FMs, and preserving their values is crucial for maintaining quantized model accuracy. Importantly, from the green box-plots, we observe that modern FMs on average possess $>0.5\%$ adjacent outliers per layer, with some FM layers showing peaks of $>2\%$. This is in stark contrast to the models evaluated by OliVe, such as BERT (Devlin et al., 2018) and OPT (Zhang et al., 2022) which have $<0.04\%$ adjacent outliers (two orders of magnitude lower than FMs like LLaMA3 and LlaVa). This indicates that while pruning values adjacent to outliers could have been ideal for models like BERT (Devlin et al., 2018), it is sub-optimal for most modern FMs as it removes crucial outlier values, leading to higher accuracy degradation. This is evident from Figure 2(b) where OliVe has significant accuracy degradation at 4-bit quantization due to its assumption on outlier locality. Unlike OliVe, MicroScopiQ does not naively prune adjacent values; instead it leverages the Hessian information (Frantar et al., 2022) to identify the least important values to prune, ensuring outlier preservation. This directly translates to high quantized model accuracy and MicroScopiQ at 2-bit consistently outperforms OliVe across different FMs.
3.3. Outlier Precision and Data Format
The ability of group A techniques like GOBO (Zadeh et al., 2020) to achieve high quantized model accuracy even at extreme quantization levels of inliers ($<4$-bits) is due to retaining outliers at higher precision. This is particularly crucial at ultra-low bit width quantization because, if inliers and outliers are to be quantized to the same precision, there will be higher outlier quantization error due to the reduced representational range. We demonstrate this effect on the MicroScopiQ quantized FM accuracy in Table 7 wherein the quantized FM has poor performance when inliers and outliers are at 2-bits compared to outliers at 4-bits. Furthermore, evidence from recent work (Wu et al., 2023) demonstrates that FP-based formats for LLMs results in superior quantization performance compared to INTs. To validate this, we compare MX-INT v/s MX-FP inlier and outlier quantization in Table 7. Evidently using MX-FP instead of MX-INT for outliers results in better performance. This is due to the higher dynamic range of FPs, which is particularly beneficial at extreme quantization levels. In this work, we quantize outliers at a higher precision (2$\times$) compared to inliers, using MX-FP for outliers and MX-INT for inliers.
4. MicroScopiQ Quantization Methodology
We present an overview of MicroScopiQ quantization in Figure 3(a) and detail it in Algorithm 1. MicroScopiQ supports various group size granularities and any inlier and outlier (2×2$\times$2 × inliers) data precision. For simplicity, we explain with inlier and outlier precision of 2/4- and 4/8-bit and group sizes of 128 for inliers and 8 for outliers.
Figure 3. (a) Overview of the proposed MicroScopiQ quantization framework depicting methodology of inlier and outlier quantization and redistribution of outlier bits for a sample LLM weight matrix. Comparison against prior quantization frameworks (b) GOBO, and (c) OliVe.
4.1. Preliminaries
The MicroScopiQ quantization framework models the layer-wise post-training quantization of FMs by partitioning each layer into multiple rows and quantizing each row at a time. Concretely, for a given input calibration dataset $\mathbb{X}$, at layer $l$ the objective is to find a quantized set of weights $\mathbb{Q}\in\mathcal{R}^{d_{row}\times d_{col}}$ that minimizes the sum of squared errors over all rows of the layer compared to the full-precision weights $\mathbb{W}\in\mathcal{R}^{d_{row}\times d_{col}}$. This can be formulated as,
(3) $\textbf{argmin}_{\mathbb{Q}}\sum_{i=1}^{d_{row}}||\mathbb{W}_{i,:}\mathbb{X}-\mathbb{Q}_{i,:}\mathbb{X}||_{2}^{2}$
We evaluate the second order derivative of Equation 3, namely the Hessian (Hassibi et al., 1993; LeCun et al., 1989) through Taylor series expansion (Hassibi et al., 1993). Note, the Hessian is the same for all rows due to its dependence only on input data and is given as, $\mathbb{H}=2\mathbb{X}\mathbb{X}^{T}$. Its inverse is, $\mathbb{H}^{-1}=(2\mathbb{X}\mathbb{X}^{T}+\lambda\mathbb{I})^{-1}$. We leverage the Hessian information to identify weights with small “saliencies” for pruning (L17 in Algo. 1). The number of inliers to prune is determined by the number of outliers, as detailed later.
To improve quantization performance, we take inspiration from (Frantar et al., 2022) and adjust the weights of unquantized rows to minimize the net error while quantizing a particular row of weights. The associated equation (L31 of Algo. 1) for weight update using the Hessian is derived by solving the Lagrangian of Equation 3. However, updating all remaining rows each time a row is quantized incurs significant compute-overhead, making this intractable for billion-scale FMs. Therefore, as pointed out in (Frantar et al., 2022), we partition the rows into non-overlapping contiguous row-blocks (rB) of size $128$ rows and localize the updates of unquantized rows within each rB. We only update the rows outside a rB (L36 in Algo. 1) once all the elements of the current block are quantized. This minimizes the number of individual updates by grouping updates together per rB and producing an order of magnitude speedup.
Input : $\mathbb{W}\in\mathcal{R}^{d_{row}\times d_{col}}$, calibration data $\mathbb{X}$, $\mathbb{H}^{-1}=(2\mathbb{X}\mathbb{X}^{T}+\lambda\mathbb{I})^{-1}$, row block (rB), macro-block (MaB) $B_{M}$, and micro-block ($\mu$B) $B_{\mu}$
1
Output : Quantized weight $\mathbb{Q}\in\mathcal{R}^{d_{row}\times d_{col}}$, perm (Permutation list)
2 # Iterate over row blocks
3for $i=0,rB,2rB,\cdots d_{col}-rB$ do
4       for $j=i,i+1,\cdots,i+rB-1$ do
5             # Step 1.0: Divide each row into non-overlapping Macro-Blocks
6            for $\mathbb{W}_{j,MaB}\in\mathbb{W}_{j,:}$ do
7                   # Step 1.1: Separate inlier and outlier in each Macro-Block
[…]
9                  # Step 1.2: Quantize Inliers to lower precision
10                  $\mathbb{Q}^{in},I_{sf}$ = InlierQuantization($\mathbb{W}^{in}$)
11                  for $\mathbb{W}_{j,\mu B}\in\mathbb{W}_{j,MaB}$ do
12                         # Step 2.0: Count Number of Outliers in a Micro-Block
13                        n = min($B_{\mu}/2$, NumOutliers($\mathbb{W}^{out}_{\mu B}$))
14                        # Step 2.1: Initialize Inlier Index List
[…]
16                         # Step 2.2: Identify n least Important Inlier Position
17                         for n iterations do
18
19                              $p=argmin_{p\in\mathbb{W}^{in}_{\mu B}}w_{p}^{2}/[\mathbb{H}^{-1}]_{pp}$
20                              # Step 2.3: Prune least important Inliers
21                              $w_{p}\leftarrow 0$
22                              # Step 2.4: Update M with the location of $w_{p}$
23                              $M\leftarrow M+\{p\}$
24                         end for
25
26                        # Step 2.5: Quantize Outliers to higher precision
27                        $\mathbb{Q}^{out},O_{sf}$ = OutlierQuant($\mathbb{W}^{out}_{\mu B},I_{sf}$)
28                        # Step 3.0: Distribute LSB Outlier Bits to Sparse Inlier Indices
29                        perm $\leftarrow$ DistributeOutlierBits($\mathbb{Q}^{out}$, $M$)
30                   end for
31
32                  $\mathbb{Q}_{j,MaB}=\mathbb{Q}^{in}+\mathbb{Q}^{out}$
33             end for
34
35            # Step 3.1: Quantization Error
36            $\mathbb{E}_{(j-i),:}=(\mathbb{W}_{j,:}-\mathbb{Q}_{j,:})/[\mathbb{H}^{-1}]_{jj}$
37            # Step 3.2: Update weights in rB to compensate quantization error
38            $\mathbb{W}_{j:(i+rB),:}=\mathbb{W}_{j:(i+rB),:}-\mathbb{E}_{(j-i),:}\cdot\mathbb{H}^{-1}_{j:(i+rB),j}$
39       end for
40      # Step 3.3: Update remaining weights after a row block is quantized
41      $\mathbb{W}_{(i+rB):,:}=\mathbb{W}_{(i+rB):,:}-\mathbb{E}\cdot\mathbb{H}^{-1}_{(i+rB):,i:(i+rB)}$
42 end for
Algorithm 1 MicroScopiQ Quantization Framework
4.2. Inlier and Outlier Weight Quantization
In Step 1 (Figure 3(a), Algorithm 1), each row to be quantized is first divided into multiple non overlapping contiguous macro-blocks (MaBs) of size $B_{M}=128$. All inliers are quantized within a MaB and share the scale factor. Each MaB is then subdivided into multiple non-overlapping contiguous micro-blocks ($\mu$Bs) of size $B_{\mu}=8$ with sixteen $\mu$Bs forming a MaB. The outliers present in each $\mu$B shares same scale(s). As depicted in Figure 3(a), Step 1, the quantization process begins by first identifying inliers and outliers in each MaB by using the $3\sigma$ rule. A shared 8-bit power-of-two scale factor ($2^{I_{sf}}$), following Equation 1 is calculated for all inliers in a MaB and the inliers are quantized to 2-bit or 4-bit, resulting in MX-INT-(2/4)128 quantization. Interestingly, we observe that the inlier scale factor in each MaB is always a negative power of two for all FMs under consideration. We leverage this observation to reduce outlier magnitude, by multiplying all outlier values in a MaB with the inlier scale factor ($2^{I_{sf}}$) (this can also be perceived as division by $2^{-I_{sf}}$, for conformity with Equation 2). This pre-processing helps make outlier quantization easier, by pre-reducing its dynamic range before the actual outlier quantization.
Unlike inliers, outliers are quantized per $\mu$B, to reduce quantization error due to shared scaling over a larger group size (see section 7). After identifying outliers present in a $\mu$B, we compute a shared 8-bit MXScale that is calculated by concatenating the level-1 power-of-two scale factor ($2^{O^{l_{1}}_{sf}}$) and level-2 microExponent ($\mu X$). The level-1 scale factor is calculated by following Equation 1 to obtain 7 or 5-bit MSBs of MXScale depending on size of $\mu X$ being 1 or 3-bit of the LSBs–corresponds to exponent size of the FP format (depicted in Step 2 in Figure 3(a)). The outliers in a $\mu$B are scaled by ($2^{O^{l_{1}}_{sf}}$), following Equation 2 and then quantized to either e1m2/e3m4 FP-format (Rouhani et al., 2023) for $b$ = 4 or 8-bit, respectively. Post quantization of outliers, the level-two scale factor or the $\mu X$ is obtained by extracting the common exponent among all outliers in a $\mu$B. This process results in a MX-FP-$b$8,8 outlier quantization. The final outlier scale factor is $2^{O_{sf}}$ where, $O_{sf}$ is expressed as ${O_{sf}}={O^{l_{1}}_{sf}+\mu X-I_{sf}}$. The term $I_{sf}$ in the final outlier scale factor accounts for multiplication by $2^{I_{sf}}$ (or division by the inverse) during outlier pre-processing.
4.3. Outlier Value Encoding through N:M Structured Pruning
To formalize Step 2, let us assume $n$ outliers are present in a $\mu$B quantized to MX-FP-44448,8. After sharing 1-bit $\mu X$ across the $n$ outliers, each outlier has 1 sign ($s$) and 2 mantissa bits ($m_{1}m_{0}$). The inliers are 2-bit 2’s complement MX-INT and has 1-bit sign and 1-bit magnitude. Since the outlier bits have two mantissa bits with one sign, to ensure symmetric outlier distribution, we duplicate their sign bit and assign each sign to a mantissa creating and partitioning into two halves Upper,Lower of size 2-bit each mimicking inlier MX-INT structure i.e., $\{sm_{1},sm_{0}\}$ (see Figure 3(a), Step 2). While we demonstrate this for a specific example, the process can be generalized to other inlier and outlier formats bit-widths wherein, the remaining outlier bits need to be reorganized into Upper and Lower distributable halves each of size $bb$ (per element bit-budget), with $bb$ representing the operational bit-width of each PE.
In Step 3, to ensure a fixed $bb$ and data-type of a layer for aligned memory access and simple PE design, we distribute the outlier LSBs (Lower half) to the least important inlier locations that are pruned within a $\mu$B. For $n$ outliers in a $\mu$B we identify $n$ least important inlier locations to prune via the Hessian information (see Algorithm 1), forming a ($B_{\mu}$-$n$):$B_{\mu}$ structured pruning pattern  [^3] Following the structure of a standard N:M pattern (Jeong et al., 2023) where N = ($B_{\mu}$-$n$) and M = $B_{\mu}$ (Jeong et al., 2023; Sun et al., 2021) i.e., ($B_{\mu}$-$n$) non-zero values exist for every $B_{\mu}$ after pruning. To keep track of the corresponding halves of each of the outliers in a $\mu$B, we maintain a per-$\mu$B permutation list (Note: if there are no outliers in a $\mu$B, a permutation list is not stored). The permutation list for each $\mu$B that has outliers is made up of $\frac{B_{\mu}}{2}$ elements (maximum number of outliers supported per $\mu B$) with each element storing the locations of the Upper and Lower halves of outliers in a 6-bit format $\{\texttt{Upper}_{loc},\texttt{Lower}_{loc}\}$ (see Figure 3(a) Step 3). Note that, in case of outlier count $>\frac{B_{\mu}}{2}$, the mB size should be chosen to be higher to prevent pruning of outliers and causing higher accuracy degradation. While such a situation does not arise in any of the models that we evaluate with $B_{\mu}=8$, this feature only serves as a demonstration of the flexibility of MicroScopiQ for future models.
Figure 4. Integration of MicroScopiQ into a WS systolic array.
Figure 5. MicroScopiQ memory organization.
4.4. Effective Bit Width (EBW) Calculation
[…]
(4) $\texttt{EBW}_{FM}=\frac{\sum_{i=1}^{l}{(x\cdot m\cdot\texttt{EBW}_{O}+(100-x)\cdot m\cdot\texttt{EBW}_{I})}/{m}}{l}$
5. MicroScopiQ Accelerator Architecture
Figure 6. Example scheduling of a GEMM operation on a $2\times 2$ MicroScopiQ accelerator.
5.1. Architecture Overview
In Figure 4, we depict the MicroScopiQ accelerator architecture, integrated into a standard weight stationary systolic array. The accelerator is optimized for throughput and all compute units (PE array, ReCoN etc.) are internally pipelined with interleaved pipeline stages. Figure 5 illustrates MicroScopiQ’s interaction with on-chip and off-chip memories. Following prior work (Keller et al., 2023; Fang et al., 2024; Dai et al., 2021), off-chip memory is modeled as HBM2 with a 256 GB/s bandwidth. The off-chip memory layout (Guo et al., 2023; Zadeh et al., 2020) of a MicroScopiQ-quantized tensor, shown in Figure 5, consists of two sections: quantized weights and metadata. Following prior work (Guo et al., 2023, 2022; Ramachandran et al., 2024b) a two-level on-chip memory hierarchy is employed: 2MB L2 global SRAM loads data from DRAM and transfers it to MicroScopiQ buffers (section 5.2) via an Open Compute Protocol OCP-SRAM interface (Sumbul et al., 2022) with a bandwidth of 64 GB/s.
Through the numbered steps in Figure 4, we explain the computational flow of the MicroScopiQ accelerator. And in Figure 6(a-c), we demonstrate the mapping and schedule for a sample $2\times 2$ GEMM operation with inlier and outlier weights, following the same computation flow.
Assume a quantized FM with outliers distributed within a $\mu$B and corresponding metadata calculated offline.
1
One $\mu$B or multiple $\mu$Bs of weights are mapped to each PE row as typically $B_{\mu}\leq$#\## of columns in PE array (Jouppi et al., 2017). Each PE in a row either receives a weight at high-precision (e.g. 4-bit) or multiple packed weights at low-precision (e.g. two 2-bits).
2
PE row 0 receives quantized input activation (iAct) from left and input accumulation (iAcc) from top. The PEs in row 0 perform multiplication of the stationary weights with the iActs, regardless of outlier/inlier weights. During accumulation, assuming the $\mu$Bs mapped to PE row 0 do not contain outliers (see Figure 6(c)), the PEs accumulate the computed multiplication result with iAcc and direct the partial sums to PE row 1.
3
PE row 1 receives partial sums from the top and similarly performs INT multiplication between weights and iAct. As shown in the scheduling figure, in parallel, the output scale factors can be calculated (section 5.5).
4
The mapped $\mu$Bs to PE row 1 have outliers, the controller directs all PE outputs to ReCoN. PE row 1 offloads the accumulation to ReCoN (shown by dotted lines in Figure 6(c)). This is needed because the outlier will have its two halves distributed in different locations and mapped to different PEs. While the PEs in row 0 with inlier weights can accumulate the partial sum output, the PEs in row 1 with the Upper and Lower outlier halves cannot compute the outlier partial sum output. After receiving outputs from row 1, ReCoN reorders and calculates FP-outlier partial sum.
5
The controller signals PE row 2 to expect input partial sums from ReCoN and not from the previous row. If PE row 1 is the last row as in Figure 6(c), the oActs are directed to the oAct buffer.
6
oActs are post-processed (scaled and quantized). The post-processing of oActs does not require all the oActs to be computed. The post processing operations can be conducted as and when the oActs are generated and overlapped with the rest of the computation so as to hide computation and memory access latency overhead.
5.2. On-chip Storage and Control
Weight/Activation Organization. The input activations (iActs) are stored in the iAct buffer as 8-bit INTs. Lower-precision iActs ($<$8-bit) are supported through sign-extension. The same process is followed for oActs (output activations). The weight buffer stores weights at 4-bit granularity. At lower precision (2-bits), each buffer location simultaneously stores two 2-bit weights. The MODE signal from the controller delineates the bit format of weights in the weight buffer i.e, one 4-bit weight or two 2-bit weights.
Instruction Buffer (IB). It stores all the metadata required for inference of MicroScopiQ quantized FMs. Particularly, the IB stores the outlier distribution permutation list i.e., configurations for ReCoN and scale factors.
MicroScopiQ Controller. The controller generates appropriate control and signals to all units. It is functionally very similar to standard systolic array controllers (Sharma et al., 2018), however, with added functionality to support specific features of ReCoN (arbitration of simultaneous access by multiple rows), multi-precision PEs, post-processing unit. Furthermore, unlike traditional controllers, the MicroScopiQ controller exerts fine-grained control using handshaking signals on the streaming of iActs, iAccs into the PE array to account for the increased pipeline depth through ReCoN for rows with outliers. The controller also generates OCP-SRAM interface control signals to manage data flow between the buffers and the L2 global SRAM.
5.3. Multi-Precision PE Array
Multi-Precision. We depict the multi-precision MicroScopiQ PE in Figure 7(a). Existing multi-precision accelerators (Guo et al., 2023; Sharma et al., 2018; Guo et al., 2022) typically follow a bottom-up approach, employing low-precision PEs and grouping neighboring PEs to support higher precision. This sacrifices throughput for multi-precision support, as multiple PE columns are required to perform a single MAC operation. This reduces parallelism and increases latency. We adopt a different strategy via the MODE signal for multi-precision support by mapping multiple weights that share the same rightbound iAct to the same PE at lower precision (2-bits) or a single weight at higher precision (4-bits). The two 2-bit weights mapped to the same PE in MicroScopiQ are the weights that would have been mapped to different columns within the same row in existing multi-precision accelerators. At high-precision operation we utilize available parallelism from the PE array while at lower-precision we increase throughput by the parallel evaluation of multiple partial sums.
Multiplication Stage. Inspired by (Liu et al., 2023; Umuroglu et al., 2018) we present a multiplier-tree architecture for multi-precision INT multiplication. The weights and iAct are partitioned across the four $4$-bit $\times$ $2$-bit multipliers to calculate partial sums ($P_{00},P_{01},P_{10},P_{11}$) as described in Figure 7(a). Based on the MODE signal the partial sums above are combined using a combination of adders and shifters to provide the result (Res) of different bit-precision weights with iAct as follows,
(5) $\small\texttt{Res}=\begin{cases}\{(P_{11}<<2+P_{10}),(P_{01}<<2+P_{00})\},\textbf{MODE}_{2b}\\ (P_{11}<<4+P_{00})+(P_{01}<<2+P_{10}<<2),\textbf{MODE}_{4b}\end{cases}$
Figure 7. Microarchitecture of (a) Multi-precision PE, (b) Synchronization buffer and c) ReCoN switch.
Accumulation Stage. It receives the MUL-stage output (Res) and also iAcc from either ReCoN or the previous PE row. The controller signals the ADD stage via Outlier_Present to modify accumulation behavior based on whether inlier/outlier weights are present in the PE. If the Res corresponds to a multiplication of inlier weight with iAct, then the corresponding Res is added with iAcc using two adders with a multiplexer that enables multi-precision by propagating the carry from one adder to other. In low precision MODE, the two adders work independently to calculate partial sums in parallel. For high precision MODE, both adders work together to produce a single accumulation result with carry propagation through the multiplexer. On the other hand, if Res corresponds to multiplication of either of the outlier halves with iAct, the actual outlier accumulation is offloaded to ReCoN by concatenating Res and iAcc444the concatenated output is notated as $O_{\texttt{Upper},\texttt{Res}}$ or $O_{\texttt{Lower},\texttt{Res}}$.. This is done to prevent incorrect outlier partial sum calculation. The controller directs partial sum outputs to ReCoN or next PE row based on presence of outliers, via the OAcc_NoC/PE signal.
5.4. Redistribution and Coordination NoC
The Redistribution and Coordination NoC (ReCoN) is a multistage butterfly NoC in the MicroScopiQ accelerator time-multiplexed and shared across PE rows (see section 7.8). This is a more cost effective way of handling outliers compared to OliVe and GOBO (Guo et al., 2023; Zadeh et al., 2020) that handle outliers within the PE. Since PEs are large in number OliVe, GOBO will incur larger costs compared to MicroScopiQ.
ReCoN topology. ReCoN is composed of $n(\log_{2}(n)+1)$2-input, 2-output} ReCoN switches, in a multistage butterfly NoC topology (Krishna et al., 2014). The input and output stages of ReCoN also employ {2-input, 2-output} switches, with a dedicated switch for each column receiving (transmitting) partial sums into (from) one of the two input (output) ports. The other port of the input and output stages of each switch is tied to 00. Each ReCoN stage also receives the same iActs used by the PEs to compute the partials sums. As we shall show later, this is done to facilitate FP-outlier’s hidden-bit processing.
Figure 8. End-to-end example illustrating the working of MicroScopiQ with a $4\times 4$ PE array (only two rows shown).
ReCoN Input Interface. ReCoN is shared and time-multiplexed across all PE rows, and at a particular instant many rows might require access to it. At the inputs, the column-wise arbiters guided by control signals from the controller, coordinate and resolve contention between different PE rows and ensure a fair scheduling (for simplicity in Figure 4, we conceptually depict this operation as an N-input multiplexer). Due to the skewed data flow of systolic array computations (Jouppi et al., 2017; Ramachandran et al., 2024b), partial sums from different columns of the same row do not arrive at the ReCoN input at the same cycle. A synchronization buffer (Figure 7(c)) is positioned between ReCoN and column-wise multiplexers which propagates the partial sums through sets of buffers of different lengths (column producing the partial sum fastest has the most buffers), synchronizing the arrival of inputs from different columns. The synchronization buffer sends an ACK signal to the PE row whose values have been accepted so that it can progress in its computation and the other simultaneously contesting rows will be on hold and acknowledged in the next cycle. When there are N simultaneously contesting rows, the row that is acknowledged the last adds an N-1 cycle latency to its processing. As we show in Figure 14(b), Figure 18(a) the num. of conflicts in accessing ReCoN are minimal and if latency is a critical factor multiple ReCoN units can be used to minimize conflicts.
ReCoN switch. It is a {2-input, 2-output} switch (Figure 7(c)) that performs three major operations based on a 3-bit configuration:
  • Pass ($=$): Passes the input from left (right) port to left (right) output port.
  • Swap ($\times$): Directs the input from the left (right) input to the right (left) output, with the opposite output port receiving the right (left) port input, 0 or iAcc (seesection 5.6).
  • Merge (||||| |): This function is triggered when a switch receives $O_{\texttt{Upper},\texttt{Res}}$ and $O_{\texttt{Lower},\texttt{Res}}$ at its left and right ports, respectively. The switch separates the Res from iAcc (denoted as // in Figure 4) from both inputs and selects the Upper result’s iAcc, since it is the correct iAcc for accumulation ($O_{\texttt{Lower},\texttt{Res}}$ iAcc is selected during Swap). The Upper and Lower halves of outlier weight’s magnitude bits are actually mantissa bits ($<1$) of MX-FP which the PE treat as INTs. Therefore the Res of $O_{\texttt{Upper},\texttt{Res}}$ and $O_{\texttt{Lower},\texttt{Res}}$ are shifted 1×1$\times$1 × and 2×2$\times$2 × respectively (not shown in figure, but internally handles multi-precision MODE) to account for the decimal position of mantissa in the MX-FP format. The shifted Res are then accumulated with the corresponding iAcc. FP-formats also have a hidden-bit ($1.0$) (Wang et al., 2018; Ramachandran et al., 2024b); to account for the contribution of the hidden bit to the outlier partial sum we also add iAct to the accumulation above.
ReCoN Output Interface. The reordered partial sums are then routed to the subsequent row in the PE array or if the input is from the last PE row, to the oAct buffer. Since ReCoN is pipelined internally and with the rest of the PE array, reordered and processed partial sums are produced every cycle once the pipeline depth (=number of ReCoN stages) is filled.
5.5. Post Processing and Scale Compute Unit
The shared output scale factor for $B_{M}$=128 is calculated as ${oAct_{sf}}={O_{sf}+iAct_{sf}}$ (See expansion of $O_{sf}$ in section 4.2.). The calculation of $oAct_{sf}$ (with simple adders and subtractors) is independent of the processing done in the PE array. Therefore, we overlap the computation of $oAct_{sf}$ with the processing of oActs to efficiently hide the computation latency similar to prior works (Dai et al., 2021; Ramachandran et al., 2024b). Since we maintain different scales for inliers and outliers (see section 4.2), and the oAct scale factor is calculated based on the outlier scale factor ($O_{sf}$) (we found this to result in least quantization error of oActs), the oActs which are generated through computation with only inlier weights (identified through the in/out control signal) are shifted by the shift value depicted in the “Post Processing and Scale Compute Unit” in Figure 4 to ensure conformity with the final scale factor. Finally, the oActs are scaled with the computed output scale factor through a simple right shift (since it is a power-of-two scale factor, division can be implemented through right shift operation) and quantized to MX-INT-(4/8)128 before being sent to external memory or routed back to the iAct buffer for computation with the next layer’s weights. Similar to prior works (Ramachandran et al., 2024b), (Keller et al., 2023) the post processing unit is also responsible for handling all non-linear operations.
Figure 9. Illustration of W4A4 MicroScopiQ GEMM on a GPU.
Table 2. Quantization Results for LLMs. We report WikiText2 perplexity numbers (lower the better).
  OPT(Zhang et al., 2022) LLaMA-2(Touvron et al., 2023) LLaMA-3(Meta, 2024) Mixtral (Jiang et al., 2024) Phi-3(Abdin et al., 2024)
Method W/A 6.7B 175B 7B 13B 70B 8B 70B 8x7B 3.8B 14B
  Baseline 16/16 10.86 8.34 5.47 4.83 3.31 6.13 2.85 3.84 6.33 4.31
  OliVe (Guo et al., 2022) 4/16 12.20 9.09 11.52 9.34 7.23 10.29 5.65 6.19 8.57 7.81
GOBO (Zadeh et al., 2020) 4/16 10.97 8.71 5.79 5.03 3.45 7.11 3.53 4.22 6.64 4.78
GPTQ (Frantar et al., 2022) 4/16 11.12 9.09 6.23 5.58 4.28 8.12 3.75 4.68 7.17 5.13
AWQ (Lin et al., 2024a) 4/16 10.97 8.74 5.82 5.19 4.08 7.96 3.58 4.36 6.72 4.99
OmniQuant (Shao et al., 2023) 4/16 10.96 8.72 5.74 5.02 3.47 7.09 3.46 4.19 6.67 4.82
MicroScopiQ (Ours) 4/16 10.91 8.62 5.65 5.02 3.42 6.89 3.25 4.07 6.61 4.70
  OliVe (Guo et al., 2023) 4/4 55.44 14.17 19.28 14.96 13.59 27.65 9.34 23.53 17.63 15.29
OmniQuant (Shao et al., 2023) 4/4 11.61 9.88 11.47 8.32 5.41 10.21 5.30 5.98 8.21 6.40
SmoothQuant (Xiao et al., 2023) 4/4 19.54 17.62 20.47 15.63 17.62 29.54 19.32 37.54 18.11 15.39
Atom (Zhao et al., 2024) 4/4 11.15 9.02 6.16 6.12 5.20 8.12 4.69 5.35 7.59 5.95
MicroScopiQ (Ours) 4/4 10.97 8.95 6.11 5.57 4.48 8.12 4.65 5.03 6.95 5.41
  OmniQuant (Shao et al., 2023) 2/16 11.61 9.66 9.62 7.56 6.11 9.13 6.17 6.02 7.09 6.28
SDQ (Jeong et al., 2024) 2/16 12.09 10.04 10.47 8.09 6.98 10.54 6.93 7.62 7.39 6.92
MicroScopiQ (Ours) 2/16 11.51 9.42 8.43 7.06 6.01 8.97 5.91 6.02 7.16 6.03
  OmniQuant (Shao et al., 2023) 2/8 11.99 10.23 9.62 8.92 6.83 9.39 6.59 6.29 7.95 7.37
Atom (Zhao et al., 2024) 2/8 11.95 10.13 9.23 8.54 6.33 9.13 6.35 6.14 7.46 7.29
MicroScopiQ (Ours) 2/8 11.77 9.98 9.06 8.06 6.33 9.08 6.02 6.17 7.38 6.82
5.6. MicroScopiQ Walkthrough Example
In this section, we discuss an end-to-end example that concretely showcases the working of MicroScopiQ. In the example shown in Figure 8, we employ a $4\times 4$ systolic array (2 rows shown) with PE row 0 having one outlier and three inliers in the $\mu$B and PE row 1 having all inlier weights. Note: Not all rows require access to ReCoN, only rows that have one or more outlier weights (row 0) in the PE require access. We assume $\mu$B=4, inlier and outlier quantization formats of MX-INT-2128 and MX-FP-44,4, respectively.
1
To demonstrate that outlier partial sums processed through ReCoN results in accurate FP-outlier partial sum, we first show a $\mu B$ with one outlier and its actual FP-value after sharing of $\mu X$, and having a value of $1.5_{10}$. If there were no outlier distribution, the actual partial sum from an “outlier-specific PE” for the shown iAct=$32_{10}$ and iAcc=$8_{10}$ will be $56_{10}$. However as described in section 4.3, we distribute outliers within a $\mu$B into Upper and Lower halves. Note: The hidden-bit is implicit and is only shown for demonstration.
2
The $\mu$B after distributing the outlier is mapped to PE row 0.
3
PE row 0 computes partial sum outputs for inliers in columns 1 and 2, while it outputs $O_{\texttt{Upper},\texttt{Res}}$ and $O_{\texttt{Lower},\texttt{Res}}$ for the other two columns that have outliers. The output of PE row 0 is routed through ReCoN. In ReCoN, the inliers are passed down at all stages, since they are already present in their respective columns.
4
The level 1 switch at column 3, executes a swap operation and redirects the lower half outlier output towards the upper half while simultaneously sending iAcc through the other port. This iAcc corresponds to the iAcc input to column 3 PE in row 0. This is functionally correct because, the lower half of outlier is distributed in this column by zeroing out/pruning the inlier at that location, hence the partial sum output of that PE to the corresponding PE in the next row should only be iAcc since the original weight at that location is zero. Similarly in level 2 of ReCoN, the switch that receives the $O_{\texttt{Lower},\texttt{Res}}$ from level 1 again executes a swap operation. Finally in level 3 the outlier output halves are merged together to realize $O^{0}_{Res}$, while all other switches execute pass. This merge operation results in a $O^{0}_{Res}$ that matches the expected output.
5
, 6 The partial sums from ReCoN are then sent to PE row 1 which then perform MAC operations and sends it to the subsequent row since there are no outlier weights. Similarly all rows execute operations until the final partial sum is calculated.
6. MicroScopiQ Integration in GPUs
In this section, we demonstrate the integration of MicroScopiQ into GPUs via SW (section 6.1) and HW (section 6.2) modifications. For a matrix multiplication problem of size $M\times N\times K$, each GPU thread block (Lin et al., 2024b; Zhao et al., 2024) is responsible for computing a $T_{m}\times T_{n}$ output tile. This computation is performed iteratively along the $K$ dimension (Figure 9), which is referred to as the main loop. MicroScopiQ quantized models face acceleration challenges on GPUs due to: (a) high-cost pointer arithmetic from distributed outliers and (b) tensor cores’ inability to co-issue INT and FP operations (Corporation, 2017), preventing direct acceleration of tiles. Our SW, HW optimizations address these issues to maximize GPU performance.
6.1. SW: Kernel-Level Optimizations
Register Caching. We implement a virtual caching layer (Ben-Sasson et al., 2016) to efficiently distribute outliers within a warp. Instead of each thread block loading and merging outliers from shared memory, we use shfl_sync(m, r, t) (Ben-Sasson et al., 2016) for intra-warp communication. This primitive allows a thread to share its register r while reading the value from thread t within the same warp, using m as the thread selector mask. With $B_{\mu}=8$, each warp (32 threads) has 4 $\mu B$s, allowing efficient intra-warp outlier merging based on the permutation list.
GEMM on Tensor Cores. Due to MicroScopiQ’s MX-FP format for outliers and MX-INT for inliers, mixed tiles after outlier merging (iteration 0 in Figure 9) must first be dequantized to FP16 before GEMM execution. In contrast, inlier-only tiles (iteration 1) can leverage 4-bit INT tensor cores for efficient acceleration but still require INT32-to-FP16 dequantization for accumulation along the K-dimension (Figure 9) with other tiles. We implement this block-level dynamic decision to deliver maximum performance. However, the need for repeated dequantization and lack of multi-precision support is a significant bottleneck on GPUs.
6.2. HW: Potential Tensor Core Modification
Following prior modeling (Guo et al., 2023; Raihan et al., 2019), each tensor core (Corporation, 2017; Abdelkhalik et al., 2022) performs 16-bit FEDPs (four-element dot products). Therefore, each tensor core at 4-bits can conduct 16EDPs. Efficiently accelerating MicroScopiQ requires simultaneous INT+FP support within the tensor core. Following the functionality of ReCoN in section 5.4, each 16EDP operation requires a variable right shifter (Inliers: $>>0$, Outlier Upper Half: $>>1$, Outlier Lower Half: $>>2$) to account for the FP mantissa for outlier products. With typical GPU die sizes (RTX 2080 Ti: 754 $mm^{2}$), adding a shifter has negligible overhead ($\sim 0.1\%$).
7. Experimental Evaluations
7.1. Experimental Setup [^5] We denote each quantization configuration by its $bb$ to enhance readability and facilitate clear comparisons.
Models and Datasets. We evaluate on OPT (Zhang et al., 2022), LLaMA2 (Touvron et al., 2023) LLaMA3 (Meta, 2024), Phi-3 (SLM) (Abdin et al., 2024) and Mixtral (MoE) (Jiang et al., 2024) LLM model families. The VLMs evaluated are OpenFlamingo (Awadalla et al., 2023), VILA (Lin et al., 2024c) and LlaVa (Liu et al., 2024a). We use a calibration dataset consisting of 256 random samples from the PILE dataset (Gao et al., 2020), so as to not overfit to any particular downstream domain (Lin et al., 2024a). We compare different LLMs based on perplexity (PPL $\downarrow$) on the WikiText2 (Merity et al., 2016) dataset and benchmark accuracy on 6 different tasks BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-c (Clark et al., 2018), MMLU (Hendrycks et al., 2020) and Winogrande (Sakaguchi et al., 2021).
Similarly, we evaluate VLMs on 5 vision-language benchmarks: COCO captioning (Chen et al., 2015), VQAv2 (Goyal et al., 2017), VizWiz (Gurari et al., 2018), TextVQA (Singh et al., 2019), and GQA (Hudson and Manning, 2019). Additionally, to demonstrate generalizability we also evaluate Convolutional Neural Network (CNN) baselines-ResNet50, VGG16 and State Space Model (SSM) baselines-VMamba (Liu et al., 2024b), Vim (Zhu et al., 2024). For these models we employ a calibration dataset of 64 samples from the Imagenet training set (Ramachandran et al., 2025b).
Algorithm Implementation. We implement the MicroScopiQ Quantization Framework in PyTorch (Paszke et al., 2019). All FMs are quantized using a single NVIDIA H100 GPU. The complete quantization process’ runtime ranges between 30 mins.–9 hours depending on model size (3.8B to 175B), which is at par with recent SoTA techniques (Frantar et al., 2022; Zhao et al., 2024).
Algorithm Baselines. We compare the performance of our MicroScopiQ quantization framework against existing SoTA quantization algorithms: OmniQuant (Shao et al., 2023), AWQ (Lin et al., 2024a), GPTQ (Frantar et al., 2022), Atom (Zhao et al., 2024), SDQ (Jeong et al., 2024) and co-design techniques: OliVe (Guo et al., 2023), GOBO (Zadeh et al., 2020).
Accelerator Implementation. The MicroScopiQ accelerator is implemented in Verilog RTL. We perform synthesis and PnR using Synopsys Design Compiler and Innovus with a TSMC 7nm technology library (See floorplan in Figure 16(a)) for area, power and latency calculations. All MicroscopiQ design variants attains a peak clock frequency of 1 GHz. We use CACTI (Muralimanohar et al., 2009) to estimate the area and power of on-chip memories. For end-to-end performance evaluation, we design a cycle accurate simulator based on DnnWeaver (Sharma et al., 2016) and BitFusion (Sharma et al., 2018), following previous works (Guo et al., 2023; Ramachandran et al., 2024b).
Accelerator Baselines. We compare the MicroScopiQ accelerator against OliVe (Guo et al., 2023), ANT (Guo et al., 2022), GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018) and AdaptivFloat (Tambe et al., 2020) across area, power and performance metrics. To demonstrate the minimal overhead of ReCoN on NoC-based real-world industrial and academic accelerators, we also implement and compare against baseline MTIA (Firoozshahian et al., 2023) and Eyeriss-v2 (Chen et al., 2019). For a fair comparison, we ensure all designs attain a clock frequency of 1GHz and have same memory hierarchy and bandwidth. DeepScale (Sarangi and Baas, 2021) is employed to scale all designs to the 7nm process.
GPU Implementation. The optimized MicroScopiQ kernel is implemented in CUDA (Zhao et al., 2024) with a PyTorch frontend. We extend GPGPU-Sim (Bakhoda et al., 2009) and AccelSim (Khairy et al., 2018) for Ampere GPUs to integrate and evaluate the modified tensor core. Energy estimation is performed using AccelWattch (Khairy et al., 2020), GPUWattch (Leng et al., 2013). Real GPU inference throughput follows the setup in (Zhao et al., 2024), while simulated GPU results follow (Guo et al., 2023).
GPU Baselines. We compared system-level performance with FP16 TensorRT-LLM (NVIDIA, 2024) baseline and W4A4 Atom kernel (Zhao et al., 2024).
Table 3. Comparison of quantization performance for LLaMA2-70B on LLM benchmarks.
  Method W/A ARC-c HellaSwag MMLU WinoGrande
Baseline 16/16 60.50 84.30 68.90 80.60
  OliVe (Guo et al., 2023) 2/16 38.60 55.30 39.80 60.70
OmniQuant (Shao et al., 2023) 2/16 49.70 77.80 58.20 74.20
MicroScopiQ (Ours) 2/16 53.30 81.60 63.70 77.80
7.2. LLM Quantization Results
Weight-Only Quantization. In Table 2, we compare the WikiText2 PPL of different LLMs for W4A16 and W2A16 settings. MicroScopiQ consistently outperforms all the baselines across different models and quantization settings. At W4A16, MicroScopiQ achieves near-lossless quantization performance. Particularly at W2A16, the benefits of the MicroScopiQ method is evident, achieving up to a 2.04 decrease in PPL score compared to the baselines. This is due to MicroScopiQ’s ability to quantize outliers at higher-precision and in FP-format to reduce quantization error. At W2A16, techniques like AWQ and OliVe have unacceptable accuracy degradations ($\geq 1e4$ PPL, not depicted in table). MicroScopiQ outperforms the contemporary SDQ (Jeong et al., 2023) at W2A16 with up to a 2.06 lower PPL score, owing to its flexible pruning+quantization strategy. In contrast, SDQ relies on a rigid N:M pattern, limiting its adaptability across model families. MicroScopiQ also outperforms the SoTA quantization framework OmniQuant across all models which requires 56×5-6$\times$5 - 6 × higher runtime than MicroScopiQ.
Figure 10. Weight-only quantization for VLMs, OpenFlamingo-9B (Awadalla et al., 2023) on COCO, VQAv2 and VILA-7B (Lin et al., 2024c) on VizWiz and TextVQA.
Figure 11. Qualitative results of OpenFlamingo-9B (Awadalla et al., 2023) on 8-shot COCO captioning with weight-only quantization using OliVe (W4) and MicroScopiQ (W2). Text in red and green indicate incorrect and accurate captioning respectively.
Weight-Activation Quantization. Similarly, we compare the quantization performance of MicroScopiQ for W4A4 and W2A8 configurations in Table 2. Due to the dynamic nature of activations, very rarely do techniques quantize outliers in activations directly (Zhao et al., 2024). Instead, most techniques (Xiao et al., 2023; Shao et al., 2023) migrate the activation outliers to weights to enable simpler activation quantization. We borrow the migration strength ($\alpha$) hyper-parameter introduced in SmoothQuant (Xiao et al., 2023) to migrate the difficulty of quantizing outliers in activations to weights. We find that prior techniques (Xiao et al., 2023; Shao et al., 2023) can only migrate up to 50% of activation quantization difficulty to weights, i.e., $\alpha=0.5$ before beginning to introduce higher weight quantization error. MicroScopiQ’s robustness to higher presence of outliers in weights by effectively quantizing outliers at higher precision and identifying least important weights to prune, allows $\alpha$ as high as 0.7 i.e., migrating most of the activation outliers to weights. This enables simpler activation quantization with MX-INT-(4/8)128 and absolving the need to handle outliers in activations. With $\alpha=0.7$ for MicroScopiQ, $\alpha=0.5$ for SmoothQuant and learned migration strength for OmniQuant in Table 2, MicroScopiQ outperforms all baselines with up to 7.4$\times$ lower PPL score across quantization settings. Furthermore, MicroScopiQ outperforms a recent work Atom (Zhao et al., 2024) that performs activation quantization with up to a 0.33 lower PPL score.
Benchmark Accuracy. In Figure 2(b), we compare the zero-shot task accuracy of W2A16 MicroScopiQ against W416 OliVe for three different LLM benchmarks. MicroScopiQ at lower-precision than OliVe achieves $\geq$8higher accuracy on the benchmarks compared to OliVe. OliVe shows poor accuracy due to its assumption on outlier locality, resulting in unintended outlier pruning. In Table 3 we also compare MicroScopiQ against OliVe and OmniQuant at W2A16 setting on 4 other LLM benchmarks. MicroScopiQ consistently outperforms all baselines by up to 9% across all benchmarks.
Effective Bit Width. For $bb=2,4$, MicroScopiQ has an EBW of $2.36$ and $4.15$ bits respectively. Techniques like GOBO, have a very high EBW of $15.6$, $18.17$ bits. All other baselines employ software-managed metadata and have an EBW=$bb$.
Table 4. Quantization results for non-transformer models. We report Top-1 accuracy (%) on ImageNet (higher is better).
  Method W/A ResNet50 VGG16 VMamba-S Vim-S
  Baseline 16/16 76.15% 71.59% 83.60% 80.50%
  HAWQ (Dong et al., 2019) 2/4 73.17% 68.81% - -
QMamba (Li et al., 2025) 4/4 - - 40.36% 68.50%
MicroScopiQ 4/4 75.08% 70.84% 70.07% 71.52%
MicroScopiQ 2/8 75.12% 70.87% 66.52% 71.98%
MicroScopiQ 2/4 73.61% 69.12% - -
Figure 12. Iso-accuracy comparison of different accelerators. MicroScopiQ-v1 and -v2 are executed on the same accelerator and correspond to two different data-precision distributions.
7.3. VLM Quantization Results
In Figure 10, we compare 0-shot and multi-shot (4,8, 16, 32) weight-only quantization performance of OpenFlamingo-9B on the COCO captioning task and VILA-7B on VizWiZ, TextVQA benchmarks. At W4A16 quantization, MicroScopiQ consistently outperforms all baselines and achieves on average less than 1% accuracy drop compared to the full-precision baseline, demonstrating the flexibility and widespread applicability of our method. Furthermore, at W2A16 quantization, MicroScopiQ achieves high accuracy ($<4\%$ accuracy drop), outperforming several W4A16 baselines.
We also quantitatively evaluate the performance of MicroScopiQ W2A16 quantized OpenFlamingo-9B VLM on the quality of generated captions for images drawn from the COCO captioning task in Figure 11. The MicroScopiQ quantized model generates accurate captions and preserves overall word semantics compared to OliVe. For instance, in the second figure, OliVe mislabels boat as a van, whereas MicroScopiQ is able to correctly identify it as a boat.
7.4. CNN and SSM Quantization Results
In Table 4, we demonstrate MicroScopiQ quantization on CNNs and SSMs. MicroScopiQ achieves near lossless performance at W4A4, W4A8 configurations and $\leq 3\%$ accuracy drop at W2A4 setting. Similarly, MicroScopiQ achieves up to $30\%$ higher accuracy compared to the SoTA, QMamba (Li et al., 2025), across SSM models.
7.5. Accelerator Results
Compute Area. In Table 5, we compare the accelerator compute area breakdown of MicroScopiQ (w/ 1 ReCoN unit) with baselines GOBO and OliVe for a $64\times 64$ array design. For a fair evaluation, all accelerators have identical configurations with same number of PEs. MicroScopiQ has a very low compute overhead $8.63\%$, compared to OliVe’s $9.90\%$. This is because, units like ReCoN which perform three simple operations for outlier processing has a very low area utilization compared to OliVe’s encoders and decoders. ReCoN is also time-multiplexed across all rows resulting in minimal area overhead (see section 7.7). Furthermore. the simple INT operations of MicroScopiQ allows the packing of more compute power with minimal overheads. A similar multi-precision design in OliVe would require 3.5-4$\times$ higher multi-precision support area due to complex exponent-integer PEs. GOBO has higher compute area compared to OliVe and MicroScopiQ due to large per-PE area.
Table 5. Compute area and density comparison for a $64\times 64$ array at 7nm process technology.
  Architecture Component $\langle$ Area ($\mu m^{2}$), # Units $\rangle$ Compute area ($mm^{2}$) Compute overhead Compute density (TOPS/$mm^{2}$)
[…]
Maximum Performance Per Unit Area ($TOPS/mm^{2}$). We also estimate the peak compute throughput per unit area of each accelerator (compute density), using LLaMA3-8B as the workload in Table 5. To achieve peak throughput for MicroScopiQ, FMs must be quantized at $bb=2$. MicroScopiQ achieves nearly 2$\times$ and 14$\times$ higher compute density than OliVe and GOBO respectively.
[…]
Iso-Accuracy Energy Comparison. In Figure 12(c), we show the normalized energy consumption of different accelerators, composed of static and dynamic energy. MicroScopiQ v2 has the lowest energy consumption. Compared to designs like GOBO, OLAccel and AdaptivFloat, where computations happen at higher-precisions in 8-bit and 32-bit PEs, MicroScopiQ v2 has a significant advantage in terms of the core and DRAM dynamic energy consumption with simple INT PEs. On average MicroScopiQ v2, has 1.5$\times$ lower energy consumption compared to baselines across different FMs.
Power Breakdown. The power breakdown of MicroScopiQ varies depending on the outlier distribution. For a LLaMA2-7B, the PE array consumes 56.23% of power, followed by on-chip memory (36.80%) and ReCoN (5.94%). Conversely, VILA-7B, characterized by a higher outlier percentage, exhibits increased power consumption in the ReCoN unit (7.65%) and slightly reduced on-chip memory (35.32%), PE array (55.98%) power consumption.
Figure 13. Comparison of MicroScopiQ accelerators with A100 GPU: (a) Normalized latency, (b) Normalized energy.
7.6. GPU Results
Throughput on real GPU. In Table 6, the unoptimized W4A4 MicroScopiQ (MS no-optim) underperforms the FP16 baseline due to the overhead of outlier merging in shared memory and GEMM execution in FP16. In contrast, our optimized W4A4 kernel (MS optim.) (section 6.1), achieves similar performance to SoTA technique Atom (Zhao et al., 2024) due to efficient register caching and dynamic GEMM acceleration, while simultaneously possessing better accuracy.
MicroScopiQ GPU simulation. As shown in our simulated A100 GPU results (Table 6), our tensor core modification enables low-precision INT+FP GEMM, eliminating costly dequantization and FP16 GEMMs, achieving the highest throughput.
Table 6. Normalized token generation throughput comparison across different quantization methods (Wikitext perplexity in brackets) on A100 GPU and GPU simulation (Bakhoda et al., 2009).
  Device Method LLaMA-2 13B LLaMA-3 8B
  A100 TRT-LLM FP16 1.00 (4.83) 1.00 (6.13)
W4A4 Atom 2.25 (6.12) 1.05 (8.12)
W4A4 MS no-optim. 0.98 (5.57) 0.92 (8.12)
W4A4 MS optim. 2.06 (5.57) 1.01 (8.12)
  GPGPU Sim W4A4 MS w/ New MTC 4.31 (5.57) 1.78 (8.12)
      MS: MicroScopiQ; MTC: Modified Tensor Core
GPU v/s MicroScopiQ Accelerator. We compare a standard A100 GPU with the MicroScopiQ accelerator under iso-bandwidth (off-chip: 2 TB/s, on-chip: 1.5kb/cycle/warp) (Raihan et al., 2019) and iso-compute (55,296 multipliers (NVIDIA, 2025)) scenario. As shown in Figure 13(a), MicroScopiQ v1 (W4A4) and v2 (WxA4) (see above) yield 1.2$\times$ and 1.7$\times$ speedup, respectively, over A100 (W4A4), due to multi-precision PEs, avoiding FP16 computation. Additionally, ReCoN’s reordering is pipelined with partial sum reduction, significantly outperforming GPU shfl_sync. Additionally, MicroScopiQ delivers superior energy efficiency (Figure 13(b)), whereas A100 incurs higher on-chip energy consumption, due to register-level reordering and FP16 overhead. Notably, traditional GPUs lack WmAn ($m\neq n$) acceleration, making MicroScopiQ accelerator an efficient alternative.
Table 7. Effect on Llama3-8B PPL (Meta, 2024) upon progressive inclusion of different quantization techniques.
  Quantization Method WikiText2 PPL $\downarrow$
  Baseline W16A16 6.13 (-)
$+$ Quantize all weights to INT-4 10.27 ($\uparrow 4.14$)
$+$ Quantize all weights to MX-INT-4128 9.53 ($\downarrow 0.74$)
$+$ Quantize all weights to MX-INT-2128 39.48 ($\uparrow 29.95$)
$+$ Quantize outliers to MX-FP-4128,128 10.96 ($\downarrow 28.52$)
$+$ Quantize outliers to MX-FP-48,8 8.93 ($\downarrow 2.03$)
$+$ Reduce outlier mag. by ×$\times 2^{I_{sf}}$ 8.89 ($\downarrow 0.04$)
$+$ Prune least imp. inliers per $\mu B$ 9.02 ($\uparrow 0.13$)
$+$ Compensate quantization errors/rB 8.97 ($\downarrow 0.05$)
$+$ Quantize activations to MX-INT-8128, $\alpha=0.7$ 9.08 ($\uparrow 0.11$)
$+$ 2-bit KV-cache quantization (Liu et al., 2024c) 9.58 ($\uparrow 0.50$)
Table 8. LLM quantization performance of MicroScopiQ integrated with OmniQuant.
  Method W/A Llama-2 13B Llama-3 70B Phi-3 3.8B
Baseline 16/16 4.83 2.85 6.33
  OmniQuant (Shao et al., 2023) 4/16 5.02 3.46 6.67
Omni-MicroScopiQ 4/16 4.87 2.97 6.52
  OmniQuant (Shao et al., 2023) 2/16 7.56 6.17 7.09
Omni-MicroScopiQ 2/16 6.58 5.09 6.89
  OmniQuant (Shao et al., 2023) 2/8 8.92 6.83 7.95
Omni-MicroScopiQ 2/8 7.12 5.74 7.21
7.7. Algorithm ablations
Per-component accuracy impact. In Table 7 we examine the accuracy gained or lost by progressively incorporating the quantization techniques proposed in MicroScopiQ. We initially adopt INT-4 scalar quantization on the LLaMA3-8B full-precision baseline. We then apply MX-INT-4128. However, on decreasing precision to 2-bit INTs i.e., MX-INT-2128, we observe a spike in PPL, due to higher outlier quantization error. Upon quantizing outliers to MX-FP-48,8, we regain the lost performance due to preserving outlier values at higher-precision and FP-format. Furthermore, we observe that pruning of least important weights causes a minor increase in PPL, which is regained by Hessian update. We observe that quantizing activations with simple MX-INT-8128 with a high $\alpha$ results in a very minimal increase in PPL due to the robustness of MicroScopiQ.
KV-cache quantization. In Table 7 we also report the impact of KV-cache quantization on performance. Following the 2-bit KV-cache quantization technique proposed in (Liu et al., 2024c) we quantize K per channel, and the V per token with a MaB of size 128 and residual token length (R) (Liu et al., 2024c) of 128 for both key and value.
Omni-MicroScopiQ. Our method is orthogonal to the techniques proposed in OmniQuant. We combine our method with OmniQuant to further improve the quantization performance of MicroScopiQ. We employ OmniQuant’s Learnable Weight Clipping (LWC) to learn both the inlier and outlier scale factors of MicroScopiQ and leverage the Learnable Equivalent Transformation (LET) to learn to migrate the quantization difficulty of activations to weights. In Table 8 we compare the performance of MicroScopiQ enhanced with OmniQuant (Omni-MicroScopiQ) against OmniQuant. Omni-MicroScopiQ delivers improvements up to 22%.
Figure 14. Effect of outlier group size on the Wikitext-v2 perplexity (PPL) and EBW of a MicroScopiQ quantized LLaMA3-8B. $\infty$ is per-tensor granularity.
Outlier group size. In Figure 14, we analyze the effect of different $B_{\mu}$ values on outlier diversity (measure by the standard deviation: red line), PPL (green bar), and EBW (blue line) for LLaMa3-8B. For $B_{\mu}=2,4$, accuracy degradation occurs due to outlier pruning, as many mBs have $\geq B_{\mu}/2$ outliers. As $B_{\mu}$ increases, outliers further apart show greater diversity, evidenced by the higher standard deviation. For $B_{\mu}\geq 32$, this results in increased quantization error and higher PPL, as MX-FP scales are shared across more diverse outlier values. Larger group sizes also significantly increase EBW due to metadata overhead. A balance is achieved at $B_{\mu}=8$.
Figure 15. Different variants of MicroScopiQ accelerator with varying number of ReCoN units.
7.8. Analysis of time-multiplexed ReCoN
Accelerator design variants. In Figure 15, we show the layout of different variants of MicroScopiQ with varying number of ReCoN units, each of which can be employed based on application requirements. For area critical applications, design A, with a single ReCoN can be employed to have minimal area utilization with a slight increase in latency due to access conflicts to ReCoN. With the addition of more ReCoN units, fewer rows of PEs share a single ReCoN, dramatically reducing access conflicts to the shared ReCoN unit (design B). When the number of ReCoN units equals the number of rows, each PE row gets a dedicated ReCoN unit as shown in design C, which can be employed for latency critical applications. As we shall demonstrate subsequently, we identified that in up to a $128\times 128$ PE array size, a total of 8 ReCoN units are sufficient to achieve peak performance with zero access conflicts to ReCoN.
Optimal # of ReCoN units. For a $64\times 64$ MicroScopiQ accelerator, we study the number of access conflicts for different number of ReCoN units in Figure 16(b). Evidently, the number of access conflicts to ReCoN which is measured as the percent of total number of accesses to ReCoN that result in conflicts is under 3%. With the progressive addition of more ReCoN units, the access conflicts tend to 0%. As observed in Figure 12, despite the access conflicts, 1 ReCoN unit is sufficient to reach optimal inference performance in an iso-accuracy scenario. For latency critical applications, in Figure 18(a), we quantify the performance improvement and impact on compute area by incorporating multiple ReCoN units for a LLaMA3-8B FM. Up to 8 ReCoN units enables peak performance with 21% improved latency and only 1.58×1.58$\times$1.58 × higher compute area. For large array sizes, up to 8 ReCoN units are sufficient for peak performance.
Figure 16. (a) Annotated MicroScopiQ floorplan ($64\times 64$ array). (b) Percentage access conflicts to ReCoN for $64\times 64$ PE array.
7.9. Scalability and Timing Analysis
Scalability. To study how MicroScopiQ scales to different PE array sizes, we compare the total area (including on-chip weight and iAct/oAct buffers) against OliVe at three different scales: $8\times 8,16\times 16,128\times 128$ in Figure 17. Following (Sumbul et al., 2022), for $8\times 8$ array we employ 16 kB iAct and oAct buffers and 32 kB weight buffer. We also depict the accelerator floorplan for a $64\times 64$ array in Figure 16(a). We progressively scale the buffers for larger arrays to ensure full array utilization. Since varying the number of ReCoN units trades off area for performance, we also compare three versions of MicroScopiQ with 1, 2, 8 ReCoN units. For the single ReCoN variant of MicroScopiQ, as the design scales to higher array sizes, the area overhead of ReCoN decreases ($128\times 128$ has overhead of 3%), with the on-chip area being dominated by the large PE array and buffers. We observe similar trends for 2 or 8 ReCoN units. At large PE array sizes such as $128\times 128$, adding 8 ReCoN units results in a minimal overhead of just 11%. Particularly at large array sizes the column-wise bus from PEs to ReCoN can be routed over logic as the PE array scales. With same buffer configurations, OliVe has a higher area compared to a single ReCoN MicroScopiQ variant, while comparable area to MicroScopiQ with 8 ReCoN units. At all variants, MicroScopiQ has 23×2-3$\times$2 - 3 × higher performance than OliVe (see Table 5).
Timing Analysis. MicroScopiQ at all scales attains a peak clock frequency of 1 GHz. The critical path for MicroScopiQ is from the local weights registers to multipliers within the PE. The column-wise bus (the longest wire in MicroScopiQ) from PEs to ReCoN is not on the critical path and therefore does not limit scalability.
Figure 17. Comparison of area of MicroScopiQ and OliVe at different PE array sizes. Three different versions of MicroScopiQ are depicted with different number of ReCoN units.
Implementation overhead on NoC-based accelerators. The global column-wise bus is a design commonly seen in AI ASICs like (Chen et al., 2019; Firoozshahian et al., 2023; Shao et al., 2019; Tong et al., 2024). These global buses are used in conjunction with the NoC in these designs for data distribution, layout reordering (Tong et al., 2024). We modify two large-scale accelerators MTIA and Eyeriss-v2 that employ NoCs in their design to support MicroScopiQ, by incorporating ReCoN functionality in the accelerator NoCs and modifying PEs to support all MicroScopiQ specific operations. As shown in Figure 18(b), compared to the baseline design of MTIA and Eyeriss-v2, integrating ReCoN and MicroScopiQ PEs results in only a $3\%$ and $2.3\%$ increase in compute area, respectively. Since these designs already employ NoCs, integration overhead of MicroScopiQ is minimal. This demonstrates that the benefits achieved by MicroScopiQ are attained at minimal to no additional cost.
8. Related Work
Model compression for LLMs. LLM weight compression schemes like quantization often rely on mixed-precision quantization with different outlier and inlier bit-widths (Lee et al., 2024; Dettmers et al., 2023; Zhao et al., 2024; Jeong et al., 2024; Huang et al., 2024). Other class of quantization rely on channel-wise shared scaling with different scale factors for outliers (Lee et al., 2024; Shao et al., 2023; Huang et al., 2024). Other quantization techniques mitigate outliers from weights and activations (Liu et al., 2024d; Xiao et al., 2023; Ashkboos et al., 2024b).
LLM weight pruning methods (Frantar and Alistarh, 2023; Yin et al., 2023; Ashkboos et al., 2024a; Ramachandran et al., 2025a) rely on approximate solvers to identify the importance of weights based on magnitude of weights or weights activation combined (Sun et al., 2023). However, they fail to achieve satisfactory performance even at low sparsity levels(Lu et al., 2024). MicroScopiQ provides a unified solution for outlier-aware quantization by enabling flexible pruning of least-important weights to distribute additional outlier bits, thereby, simultaneously achieving high accuracy and hardware efficiency.
Unified pruning and quantization. SDQ (Jeong et al., 2024) combines pruning and quantization by decomposing weights into inlier and outlier vectors, each employing a fixed N:M pattern and different bit precisions. In contrast, MicroScopiQ removes the rigid N:M constraint and achieves hardware efficiency through a consistent bit-budget and unified data type per element. Moreover, unlike SDQ’s sequential pruning-then-quantization approach, MicroScopiQ integrates pruning implicitly within quantization, enabling tighter coupling.
Accelerator for quantized models. Accelerators like GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018) use high-precision PEs for outliers and low-precision PEs for inliers. OliVe (Guo et al., 2023) incorporates encoding/decoding units in a PE array for inlier-outlier formats, while other accelerators (Tambe et al., 2020; Ramachandran et al., 2024b; Guo et al., 2022) propose adaptive formats and hybrid FP PEs. However, they often suffer from unaligned memory access and large overheads. MicroScopiQ accelerator introduces a novel low overhead NoC called ReCoN, that effectively abstracts FP-format complexity from INT-PEs
NoCs in ML Accelerators. NoCs are widely used in accelerators: (Qin et al., 2020; Muñoz-Martínez et al., 2023) employ a Benes NoC for data distribution, (Tong et al., 2024) uses a butterfly NoC for layout transformation, and (Firoozshahian et al., 2023; Chen et al., 2019) utilize mesh NoCs. ReCoN is a butterfly NoC employed to compute FP partial sums.
Figure 18. (a) Effect of time-multiplexed ReCoN units on the compute area and inference latency, (b) Implementation overhead of MicroScopiQ in NoC-based accelerators (Firoozshahian et al., 2023; Chen et al., 2019).
9. Conclusion
In this paper, we introduce MicroScopiQ, a novel co-design technique for post-training quantization of FMs. MicroScopiQ addresses outlier-aware quantization by: a) using higher-precision MX-FP for outliers and lower-precision INTs for inliers to maintain accuracy, and b) pruning to distribute additional outlier bits, ensuring memory alignment and hardware efficiency. The MicroScopiQ accelerator, featuring multi-precision PEs and a ReCoN NoC, efficiently handles distributed outliers. MicroScopiQ is the first PTQ technique to achieve SoTA compression for both LLMs and VLMs at just $2.36$-bits, offering 3×3$\times$3 × better inference performance and 2×2$\times$2 × lower energy consumption than existing accelerators.
Acknowledgements.
This work was supported in part by CoCoSys, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. The authors would like to thank the anonymous reviewers for their valuable feedback and suggestions, which helped improve the quality of this paper. The authors also thank Jiawei Hu for place-and-route, Jianming Tong for help with designing ReCoN, and Zishen Wan for functional accuracy simulation and constructive suggestions.